OpenStudy (anonymous):

x-9=6x+5 please help

OpenStudy (anonymous):

First, subtract x from both sides. Can you do it? :o)

OpenStudy (dmezzullo):

or u could subtract 5 from both sides

OpenStudy (anonymous):

im not sure how to do this at all

OpenStudy (dmezzullo):

ok....

OpenStudy (anonymous):

x-9=6x+5 -x -x -9=5x+5 Does that make sense?

OpenStudy (anonymous):

You're trying to get x alone. :o)

OpenStudy (anonymous):

lol i know but i am not sure how to do it im no good at math and i need it explained

OpenStudy (anonymous):

No problem. You're trying to get x alone, so, you're going to do the opposite operations to both sides. Example: + opposite is - * opposite is /

OpenStudy (anonymous):

Let me show you two steps and you try to do the third.

OpenStudy (anonymous):

x-9=6x+5 So here, we're going to get x on the same side first. By doing this, we're going to subtract x from both sides so that x will cancel out on the left: x-9=6x+5 -x -x -9=6x+5 Now, we have x on both sides, so we can start to get x alone on the right side. So, do you see how +5 is on the side with the x? So, we're going to -5 from both sides to help get x alone. So: -9=6x+5 -5 -5 -14=6x Can you tell me what we're going to do now to get that x alone? It's going to be our last step. :o)